OpenStudy (anonymous):

How many solutions exist for the system, x + y = 1 and y = -x + 1? A. no solution B. one solution C. infinitely many solutions

OpenStudy (jamesj):

Notice that the two equations are identical because y = -x + 1 <=> y + x = x - x + 1 = 1 Hence you only have one equation x + y =1. How many solutions are there to that equation?

OpenStudy (anonymous):

1?

OpenStudy (jamesj):

Well, x =1, y = 0 is one solution x = 2, y = -1 is another solution So is x = 0, y =1 In fact, if you drew the graph of that equation, there would be an infinite number of points on it, reflecting the fact there are an infinite number of answers.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (jamesj):

In general, if you have two variables you need two equations to have a unique solution. With two variables and one equation, there are an infinite number of solutions.